# How Far Are We From AGI?

### v1

@@ Abstract
The evolution of artificial intelligence (AI) has profoundly impacted human society, driving significant advancements in multiple sectors. Yet, the escalating demands on AI have highlighted the limitations of AI’s current offerings, catalyzing a movement towards Artificial General Intelligence (AGI). AGI, distinguished by its ability to execute diverse real-world tasks with efficiency and effectiveness comparable to human intelligence, reflects a paramount milestone in AI evolution. While existing works have summarized specific recent advancements of AI, they lack a comprehensive discussion of AGI’s definitions, goals, and developmental trajectories. Different from existing survey papers, this paper delves into the pivotal questions of our proximity to AGI and the strategies necessary for its realization through extensive surveys, discussions, and original perspectives. We start by articulating the requisite capability frameworks for AGI, integrating the internal, interface, and system dimensions. As the realization of AGI requires more advanced capabilities and adherence to stringent constraints, we further discuss necessary AGI alignment technologies to harmonize these factors. Notably, we emphasize the importance of approaching AGI responsibly by first defining the key levels of AGI progression, followed by the evaluation framework that situates the status-quo, and finally giving our roadmap of how to reach the pinnacle of AGI. Moreover, to give tangible insights into the ubiquitous impact of the integration of AI, we outline existing challenges and potential pathways toward AGI in multiple domains. In sum, serving as a pioneering exploration into the current state and future trajectory of AGI, this paper aims to foster a collective comprehension and catalyze broader public discussions among researchers and practitioners on AGI [^1] Project page: https://github.com/ulab-uiuc/AGI-survey. We highly welcome pull requests from the community..
@@ 1 Introduction
Figure 1: AI has Gradually Surpassed Humans. We estimate the (cumulative) percentage (in log scale) of human activities in which AI has surpassed humans in terms of competence and efficiency as we start from early embryonic systems around 1970s to more advanced ones developed recently. At each node along the polyline, we choose one representative work that revolutionized the field, bringing substantial improvements to AI technology. The trend of AI popularity and generality is increasing in an exponential rate and we can expect that starting from 2023, with the advent of work such as GPT-4 and Sora, the speed at which AI surpasses humans will increase at an unprecedented speed. The statistics presented here are calculated based on our heuristic estimations, which, despite being empirical, can serve as a rudimentary guidance for understanding the speed at which AI develops. This also provides a realistic basis for people to imagine the arrival of AGI.
@@
AGI is Within Closer Reach than Ever
As is summarized in Figure 1, the rapid development of AI has enabled its capabilities to surpass human activities in increasingly more fields, which indicates that the realization of AGI is getting closer. Hence, it is of great practical significance to revisit the question of how far we are from AGI and how can we responsibly achieve AGI by conducting a comprehensive survey that clearly establishes the expectation of future AGI and elaborates on the gap from our current AI development.
@@
Specifically, as is shown in Figure 2, we start with an overview on the major capabilities required for future AGI in terms of its internal (§ 2) competence, its connection to the external world as an interface (§ 3), and the underlying infrastructure systems (§ 4) that support all these functionalities. When it comes to deployment, more sophisticated alignment (§ 5) procedure is indispensable to unleash the growing potential of AGI systems under constraints and human expectations. Furthermore, we picture a roadmap(§ 6) where we discuss how to responsibly approach AGI, which outlines the level definition that helps locate our current state of AI, associated evaluation framework, as well as our insights to some critical problems that might hinder our progress towards AGI. Finally, we list a couple of case studies (§ 7) which concretely describe the lineage of AI technology along various domains with cautious limitations and exciting future directions. We hope that this work can lay a common ground and provide a starting point for researchers and practitioners to reflect on the state of AI and brainstorm responsible approaches to achieve AGI.
@@ 2.1 AI Perception
      Projection-based: the modality connector exists outside both the LLMs and multi-modal encoders can be quite straightforward with simple linear projections (Zhu et al., 2023a; Liu et al., 2023d; Su et al., 2023; Chen et al., 2023j; Li et al., 2024e) or incorporating relatively complex selection method (Gao et al., 2023a; Zhang et al., 2023e; Luo et al., 2023; Han et al., 2023b; Fu et al., 2024). This type of MLLMs usually activates the projection layer and/or the LLMs for two stages of alignment training.
@@
      Cross-attention-based: Flamingo (Alayrac et al., 2022) proposed the well-known perceiver with additional cross-attention mechanism inside the attention block of the LLMs. Several variants of Flamingo (Li et al., 2023k; Gong et al., 2023) also use the same or similar framework for tuning the MLLMs.
@@
      Autoregressive: MLLMs like Fuyu (Bavishi et al., 2023) and its variants (Li et al., 2023l) take vision token as the language token from the pre-training stage and use the same autoregressive training loss to update the whole model parameter.
@@
  • The diversification of modalities is essential for integrating multiple data types and improving model capabilities. It is crucial to explore less common modalities, such as graphs, and to integrate multiple modalities, such as images, audio, and video simultaneously (Han et al., 2023b; a). This will require carefully designed modules, high-quality data, and a balanced approach to managing the interplay between different modalities and their relationship with language. For example, while GPT-4V can only handle language and visual information, the recent Gemini model expands its capacity to a wider range of audio and video.
@@
  • Encouraging multi-modal systems to be more robust and reliable. As more comprehensive benchmarks covering not only general situations but also challenging inputs like math problems (Yue et al., 2023), counterfactual instructions (Zhang et al., 2023l), and attack strings (Qi et al., 2023; Zhao et al., 2023b) emerge, it becomes evident that multi-modal systems, especially the smaller ones, generally fall short in performance when facing adversarial examples and are heavily language-biased (Tu et al., 2023a; Cui et al., 2023), lacking the reasoning ability under out-of-distribution situations like multi-panel images (Fan et al., 2024), sketches (Tu et al., 2023a), long sequence images (Wang et al., 2024). These observations pose potential risks in real-world applications. To address these challenges and build more robust multi-modal AGI models, several strategies in terms of the employed learning data are considered. One approach is to incorporate adversarial examples into training (Liu et al., 2023d). Another strategy involves increasing the diversity of instruction formats (Dai et al., 2023).
@@
  • Explainable multi-modal models point out the direction for future improvement. Unlike traditional models, multi-modal models involve complex interactions between different modalities, making it essential to unravel their inner workings for understanding and creating stronger multi-modal ones. To address this, research efforts have focused on providing explanations during training or generation, offering insights into model performance and reasoning. Methods like probing model performance with diverse training data have been explored (Liu et al., 2023d; Zhao et al., 2023c; Tu et al., 2023c). Additionally, the Gemini team enhances user trust and understanding of the AI’s reasoning process by providing explanations of the generation (Team et al., 2023). Another aspect of improving multi-modal models is increasing transparency. This involves identifying the specific model components or configurations that contribute to the system’s abilities (e.g., vision encoder, connector, or training paradigms) (Wang et al., 2023d; He et al., 2024). Studies have also specifically investigated the impact of different modality processors on the overall model performance (Lin et al., 2023a; Tong et al., 2024). As multi-modal models advance, future research must prioritize explainability and transparency. This will enable us to take the full potential of these powerful AI systems while ensuring their responsible and ethical use.
@@ 2.2 AI Reasoning
Reasoning is the process of thinking logically and systematically, drawing on evidence and past experiences to form conclusions or make decisions (Fagin et al., 2004; Huang and Chang, 2022). This process is integral to daily human life and is considered a fundamental element in the development of AGI.
@@
  • Navigating through thoughts. Chain of Thought (CoT) (Wei et al., 2022b) generates a sequence of intermediate reasoning steps, known as “thoughts,” to enable models to decompose multi-step problems and allocate additional computation to more complex tasks. This offers an interpretable insight into the model’s reasoning process, helping to comprehend how an answer is derived and identify where errors in reasoning might occur. Tree of Thoughts (ToT) (Yao et al., 2023) employs tree-based search algorithms to navigate through “thoughts” for deliberate problem-solving. This allows LLMs to explore multiple reasoning paths and perform deliberate decision-making, including looking ahead or backtracking when necessary. Graph of Thoughts (GoT) (Besta et al., 2023) organizes information into a graph structure, where “thoughts” are vertices, and edges correspond to dependencies between these vertices. This graph-based organization facilitates more intricate integration and manipulation of thoughts, allowing for the creation of more sophisticated reasoning pathways and the incorporation of feedback mechanisms.
@@
  • Self-consistent reasoning. Self-consistency (Wang et al., 2022a) samples a diverse set of reasoning paths and selects the most consistent answers. This method overcomes the constraints of greedy decoding by balancing open-ended and optimal text generation, utilizing the diversity of reasoning paths to achieve more reliable outcomes. Maieutic Prompting (Jung et al., 2022) elicits explanations in an abductive and recursive manner and then deduces inferences based on these explanations and their logical relations. Progressive-Hint Prompting (Zheng et al., 2023b) employs previously generated answers as hints to progressively guide toward the correct answers, enforcing a level of self-consistency with earlier responses.
@@
  • Additional prompting strategies for enhanced reasoning. Many other prompting methods have been developed to improve the reasoning abilities of LLMs. Prompt ensembling employs a combination of prompts to generate predictions, utilizing methods such as weighted averaging (Schick and Schütze, 2020; Qin and Eisner, 2021), majority voting (Lester et al., 2021), and knowledge distillation (Schick and Schütze, 2020; Gao et al., 2020). Complexity-Based Prompting (Fu et al., 2022) creates rationales with more reasoning steps with an example selection scheme. Auto-CoT (Zhang et al., 2022b) samples questions with diversity and automatically generates reasoning chains to construct demonstrations. Least-to-Most Prompting (Zhou et al., 2022b) breaks down a complex problem into a series of simpler subproblems and then solves them in sequence. Decomposed Prompting (Khot et al., 2022) decomposes tasks into simpler sub-tasks and dynamically delegates them to sub-task-specific models.
@@
  • Dynamic reasoning and planning. ReAct (Yao et al., 2022) prompts LLMs to generate reasoning traces and action plans in an interleaved manner. This synergy between reasoning and action enables dynamic reasoning, creating, maintaining, and adjusting action plans while interacting with external environments like Wikipedia. This interaction allows the integration of additional information into the reasoning process and addresses issues like hallucination and error propagation common in chain-of-thought reasoning. “Describe, Explain, Plan and Select” (DEPS) (Wang et al., 2023b) enhances plan reliability by incorporating a dynamic feedback loop that includes description, explanation, and plan adjustment stages, significantly improving error correction and planning efficiency. Inner Monologue (Huang et al., 2022b) underscores the utility of feedback-informed planning, demonstrating improved task completion and adaptability in diverse environments by dynamically incorporating feedback to refine and adjust plans in real time. ProgPrompt (Singh et al., 2023) enables the generation of executable task plans that are both contextually relevant and adaptable to the robot’s capabilities and the environment’s state by structuring prompts as programmatic instructions and incorporating environment state feedback through assert statements. LLM+P (Liu et al., 2023b) combines the natural language processing strengths of LLMs with the precise problem-solving skills of classical planners, providing optimal solutions for planning problems that involve language description.
@@
  • Reflection and refinement. Self-refine (Madaan et al., 2024) employs iterative generation, self-generated feedback, and refinement, enabling large language models to adjust based on feedback after each generative cycle. Reflexion (Shinn et al., 2023) expands on the ReAct framework by integrating an evaluator to assess action trajectories and utilizing an LLM to generate verbal self-reflections to provide feedback for future trials. CRITIC (Gou et al., 2023) utilizes external tools, such as knowledge bases and search engines, to validate the actions produced by LLMs, then employs external knowledge for self-correction to minimize factual errors.
@@
  • Integrating language models, world models, and agent models. While language models fall short of consistent reasoning and planning in various scenarios, world and agent models can provide essential elements of human-like deliberative reasoning, including beliefs, goals, anticipation of consequences, and strategic planning. The LAW framework (Hu and Shu, 2023) suggests reasoning with world and agent models, with language models serving as the backend for implementing the system or its components. This framework combines three models in a cognitively grounded way, fostering more robust and versatile reasoning capabilities. Within this framework, Reasoning via Planning (RAP) (Hao et al., 2023) prompts an LLM to function as an agent model, guided by the same LLM acting as the world model, which predicts the next state of the reasoning after applying an action to the current state. BIP-ALM (Jin et al., 2024) uses language models as the planner in agent models, leading to an improved Theory of Mind capacity compared to using language models to infer other agents’ mental states directly. Recent studies have explored the potential for using language models to generate goals (Xie et al., 2023) or rewards (Yu et al., 2023; Kwon et al., 2023b; Ma et al., 2023a) in agent models to guide planning.
@@
  • Foundation models need to learn causation for better understanding and generalization. The foundation models rely heavily on patterns identified in their training data, which do not always capture the depth and breadth of human knowledge and experiences. Furthermore, these models often operate based on patterns extracted from data without truly comprehending the underlying causal relationships. As a result, they may encounter difficulties when presented with novel situations or those that demand a more nuanced understanding (Shi et al., 2023; Saparov et al., 2024). Future advancements in AGI should focus on learning causation over correlation, thereby achieving better generalization and deeper understanding.
@@
  • AGI should enhance social reasoning to better interact with humans and other agents. Current AI models lack a robust Theory of Mind, the ability to understand the mental states of others (Sap et al., 2022; Ullman, 2023; Jin et al., 2024). Improving this capability is essential for AGI systems to safely and effectively interact with humans and other agents in an open-ended manner. Understanding social cues and norms is central to this development, as it allows AGIs to interpret and respond to implicit communication and behavioral expectations in varied contexts (Puig et al., 2020; Zhang et al., 2023b). Advancements in social reasoning in AGI systems could lead to more empathetic and context-aware technologies, ensuring that these systems engage harmoniously and meaningfully in human societies.
  • AGI should solve the challenge in explainability and transparency, enhancing their reliability in decision-making. Current AI systems lack explainability and transparency. Operating as “black boxes,” they present challenges in understanding how they arrive at specific conclusions or answers and often lack self-explanation or provide inconsistent explanations (Chen et al., 2023k). Promising-looking techniques that elicit reasoning in natural language do not reliably correspond to the processes that the models use to reason, and the explanations generated by models can be systematically misleading (Bowman, 2023). This limitation hampers their reasoning abilities and poses significant challenges when decision-making requires justification or auditing, such as in healthcare or legal settings.
@@
  • Future AGI systems aim for dynamic reasoning across domains, ethical and efficient planning, and human-like intelligence at unparalleled scales and speeds. We are still far from achieving AGI-level capability that allows reasoning planning across varied domains without retraining or human oversight (Saparov et al., 2024). The journey involves enhancing AI systems to transfer knowledge and skills across vastly different areas, enabling them to address unforeseen situations efficiently. A key development focus is creating algorithms that can plan at various levels of abstraction, from broad strategic goals to the specifics of detailed actions. Additionally, AI systems urgently need to become better at managing resources—such as time, energy, and costs—more efficiently during the planning phases. Equally critical is ensuring that these planning processes adhere to ethical standards and safety regulations, especially in sensitive sectors, to avoid misuse or unintended outcomes. Advancements in these areas will collectively move us closer to realizing AGI with robust, versatile planning capabilities.

### v2

# How Far Are We From AGI: Are LLMs All We Need?

@@ Abstract
The evolution of artificial intelligence (AI) has profoundly impacted human society, driving significant advancements in multiple sectors. Yet, the escalating demands on AI have highlighted the limitations of AI’s current offerings, catalyzing a movement towards Artificial General Intelligence (AGI). AGI, distinguished by its ability to execute diverse real-world tasks with efficiency and effectiveness comparable to human intelligence, reflects a paramount milestone in AI evolution. While existing studies have reviewed specific advancements in AI and proposed potential paths to AGI, such as large language models (LLMs), they fall short of providing a thorough exploration of AGI’s definitions, objectives, and developmental trajectories. Unlike previous survey papers, this work goes beyond summarizing LLMs by addressing key questions about our progress toward AGI and outlining the strategies essential for its realization through comprehensive analysis, in-depth discussions, and novel insights. We start by articulating the requisite capability frameworks for AGI, integrating the internal, interface, and system dimensions. As the realization of AGI requires more advanced capabilities and adherence to stringent constraints, we further discuss necessary AGI alignment technologies to harmonize these factors. Notably, we emphasize the importance of approaching AGI responsibly by first defining the key levels of AGI progression, followed by the evaluation framework that situates the status-quo, and finally giving our roadmap of how to reach the pinnacle of AGI. Moreover, to give tangible insights into the ubiquitous impact of the integration of AI, we outline existing challenges and potential pathways toward AGI in multiple domains. In sum, serving as a pioneering exploration into the current state and future trajectory of AGI, this paper aims to foster a collective comprehension and catalyze broader public discussions among researchers and practitioners on AGI.  [^1] Project website: https://github.com/ulab-uiuc/AGI-survey. Unlike traditional publications that remain static, we embrace an innovative approach by treating this paper as a living document. We warmly welcome feedback from the community and plan to update the paper annually. Contributors on the project website will be gratefully acknowledged in future revisions.
@@ 1 Introduction
To start approaching the question of how far we are from AGI, it is important to first ground ourselves with the history of artificial intelligence advancement and understand the urge for more advanced systems. Throughout the whole paper, we hope to provide evidences and insights on where we currently stand along the road towards AGI, from the lens of many modern AI systems such as large language models. The goal is to prudently keep questioning ourselves: are LLMs all we need? It is with this enduring curiosity and awareness that we might finally begin to touch the realm of AGI.
@@
Specifically, as is shown in Figure 1, we start with an overview on the major capabilities required for future AGI in terms of its internal (§ 2) competence, its connection to the external world as an interface (§ 3), and the underlying infrastructure systems (§ 4) that support all these functionalities. When it comes to deployment, a more sophisticated alignment (§ 5) procedure is indispensable to unleash the growing potential of AGI systems under constraints and human expectations. Furthermore, we picture a roadmap(§ 6) where we discuss how to responsibly approach AGI, which outlines the three levels of AGI which are Embryonic, Superhuman, and Ultimate AGI that helps locate our current state, associated evaluation framework, as well as our insights to some critical problems that might hinder our progress towards AGI. Finally, we list a couple of case studies (§ 7) which concretely describe the lineage of AI technology along various domains with cautious limitations and exciting future directions. We hope that this work can lay a common ground and provide a starting point for researchers and practitioners to reflect on the state of AI and brainstorm responsible approaches to achieve AGI.
@@ 2.1 AI Perception
Figure 3: There are three categories for multimodal models with LLM external connections: projection-based, query-based, and language-based.
@@
      Projection-based: the modality connector exists outside both the LLMs and multi-modal encoders can be quite straightforward with simple linear projections (Zhu et al., 2023a; Liu et al., 2023d; Su et al., 2023; Chen et al., 2023j; Li et al., 2024e) or incorporating relatively complex selection method (Gao et al., 2023a; Zhang et al., 2023e; Luo et al., 2023; Han et al., 2023b; Fu et al., 2024a). This type of MLLM usually activates the projection layer and/or the LLMs for two stages of alignment training.
@@
      Cross-attention-based: Flamingo (Alayrac et al., 2022) proposed the well-known perceiver with additional cross-attention mechanism inside the attention block of the LLMs. Several variants of Flamingo (Li et al., 2023l; Gong et al., 2023) also use the same or similar framework for tuning the MLLMs.
@@
      Autoregressive: MLLMs like Fuyu (Bavishi et al., 2023) and its variants (Li et al., 2023m) take vision token as the language token from the pre-training stage and use the same autoregressive training loss to update the whole model parameter.
@@
  • The diversification of modalities is essential for integrating multiple data types and improving model capabilities. It is crucial to explore less common modalities, such as graphs, and to integrate multiple modalities, such as images, audio, and video simultaneously (Han et al., 2023b; a). This will require carefully designed modules, high-quality data, and a balanced approach to managing the interplay between different modalities and their relationship with language. For example, while GPT-4V can only handle language and visual information, the recent Gemini (Team et al., 2023) model expands its capacity to a wider range of audio and video. Potential methods for incorporating other modal perceptions: a unified modal representation tool like ImageBind (Han et al., 2023b), LangaugeBind (Zhu et al., 2023c) could bridge the modal gap and lessen the burden of learning from other modalities. Existing models that incorporate these tools have shown promising results in efficiency and task performance (Su et al., 2023; Han et al., 2023b).
@@
  • Encouraging multi-modal systems to be more robust and reliable. As more comprehensive benchmarks covering not only general situations but also challenging inputs like math problems (Yue et al., 2023), counterfactual instructions (Zhang et al., 2023m), and attack strings (Qi et al., 2023; Zhao et al., 2023b) emerge, it becomes evident that multi-modal systems, especially the smaller ones, generally fall short in performance when facing adversarial examples and are heavily language-biased (Tu et al., 2023a; Cui et al., 2023), lacking the reasoning ability under out-of-distribution situations like multi-panel images (Fan et al., 2024), sketches (Tu et al., 2023a), long sequence images (Wang et al., 2024b). These observations pose potential risks in real-world applications. To address these challenges and build more robust multi-modal AGI models, several strategies in terms of the employed learning data are considered. Future research could benefit from incorporating adversarial examples into training (Liu et al., 2023d) or involving increased diversity of training data instruction formats (Dai et al., 2023).
@@
  • Explainable multi-modal models point out the direction for future improvement. Unlike traditional models, multi-modal models involve complex interactions between different modalities, making it essential to unravel their inner workings to understand and create stronger multi-modal ones. To address this, research efforts have focused on providing explanations during training or generation, offering insights into model performance and reasoning. Methods like probing model performance with diverse training data have been explored (Liu et al., 2023d; Zhao et al., 2023c; Tu et al., 2023c). Additionally, the Gemini (Team et al., 2023) team enhances user trust and understanding of the AI’s reasoning process by providing explanations of the generation (Team et al., 2023). Another aspect of improving multi-modal models is increasing transparency. This involves identifying the specific model components or configurations that contribute to the system’s abilities (e.g., vision encoder, connector, or training paradigms) (Wang et al., 2023d; He et al., 2024). Studies have also specifically investigated the impact of different modality processors on the overall model performance (Lin et al., 2023a; Tong et al., 2024b). As multi-modal models advance, future research must prioritize explainability and transparency. This will enable us to take the full potential of these powerful AI systems while ensuring their responsible and ethical use. For example, future research avenues could explore strict controlled experiments for training AI models to decompose each part (Tong et al., 2024a) or probing model components to find the most effective module (Zhao et al., 2023c).
@@ 2.2 AI Reasoning
Reasoning is the cognitive process of drawing conclusions or making decisions based on available information, logic, and prior knowledge. It involves evaluating evidence, identifying relationships, and applying rules or principles to solve problems (Fagin et al., 2004; Huang and Chang, 2022). AI reasoning refers to the ability of AI systems to simulate this process, enabling machines to understand situations, infer conclusions, and make decisions in a way that mimics human reasoning.
@@
  • Navigating through thoughts. Chain of Thought (CoT) (Wei et al., 2022b; Kojima et al., 2022) generates a sequence of intermediate reasoning steps, known as “thoughts,” to enable models to decompose multi-step problems and allocate additional computation to more complex tasks. This offers an interpretable insight into the model’s reasoning process, helping to comprehend how an answer is derived and identify where errors in reasoning might occur. Tree of Thoughts (ToT) (Yao et al., 2023) employs tree-based search algorithms to navigate through “thoughts” for deliberate problem-solving. This allows LLMs to explore multiple reasoning paths and perform deliberate decision-making, including looking ahead or backtracking when necessary. Graph of Thoughts (GoT) (Besta et al., 2023) organizes information into a graph structure, where “thoughts” are vertices, and edges correspond to dependencies between these vertices. This graph-based organization facilitates more intricate integration and manipulation of thoughts, allowing for the creation of more sophisticated reasoning pathways and the incorporation of feedback mechanisms. Program of Thoughts (PoT) (Chen et al., 2022) leverages language models to express the reasoning process as a program, delegating the computation to an external computer that executes the generated programs to obtain the answer. This separation of computation from reasoning improves performance on problems that demand highly symbolic reasoning skills.
@@
  • Self-consistent reasoning. Self-consistency (Wang et al., 2022a) samples a diverse set of reasoning paths and selects the most consistent answers. This method overcomes the constraints of greedy decoding by balancing open-ended and optimal text generation, utilizing the diversity of reasoning paths to achieve more reliable outcomes. Maieutic Prompting (Jung et al., 2022) induces a tree of explanations abductively and recursively, then frames the inference as a satisfiability problem over these explanations and their logical relations. Progressive-Hint Prompting (Zheng et al., 2023b) employs previously generated answers as hints to progressively guide toward the correct answers, enforcing a level of self-consistency with earlier responses.
@@
  • Additional prompting strategies for enhanced reasoning. Many other prompting methods have been developed to improve the reasoning abilities of LLMs. Complexity-Based Prompting (Fu et al., 2022) creates rationales with more reasoning steps with an example selection scheme. Auto-CoT (Zhang et al., 2022b) samples questions with diversity and automatically generates reasoning chains to construct demonstrations. Least-to-Most Prompting (Zhou et al., 2022b) breaks down a complex problem into a series of simpler subproblems and then solves them in sequence. Decomposed Prompting (Khot et al., 2022) decomposes tasks into simpler sub-tasks and dynamically delegates them to sub-task-specific models. ToolLLM (Qin et al., 2023b) and ToRA (Gou et al., 2023b) integrate natural language reasoning with the use of external tools, significantly enhancing their ability to perform complex reasoning. Collaboration mechanisms between multiple agents, such as debate (Du et al., 2023), reflection (Zhang et al., 2023l), voting (Li et al., 2024g), or role-playing as different characters (Qian et al., 2023a; Zhou et al., 2023d), can further enhance their reasoning performance.
@@
  • Dynamic reasoning and planning. ReAct (Yao et al., 2022) prompts LLMs to generate reasoning traces and action plans in an interleaved manner. This synergy between reasoning and action enables dynamic reasoning, creating, maintaining, and adjusting action plans while interacting with external environments like Wikipedia. This interaction allows the integration of additional information into the reasoning process and addresses issues like hallucination and error propagation common in chain-of-thought reasoning. “Describe, Explain, Plan and Select” (DEPS) (Wang et al., 2023b) enhances plan reliability by incorporating a dynamic feedback loop that includes description, explanation, and plan adjustment stages, significantly improving error correction and planning efficiency. Inner Monologue (Huang et al., 2022b) underscores the utility of feedback-informed planning, demonstrating improved task completion and adaptability in diverse environments by dynamically incorporating feedback to refine and adjust plans in real-time. ProgPrompt (Singh et al., 2023) enables the generation of executable task plans that are both contextually relevant and adaptable to the robot’s capabilities and the environment’s state by structuring prompts as programmatic instructions and incorporating environment state feedback through assert statements. LLM+P (Liu et al., 2023b) combines the natural language processing strengths of LLMs with the precise problem-solving skills of classical planners, providing optimal solutions for planning problems that involve language description. Thought Rollback (Chen and Li, 2024) introduces a rollback mechanism that allows LLMs to revise prior steps based on error analysis, fostering adaptive reasoning by dynamically adjusting thought structures to improve problem-solving accuracy. Parameter-efficient finetuning methods (PEFTs) (Xu et al., 2023c) adapt large language models by updating only a small subset of their parameters, reducing memory usage and computational costs while achieving performance comparable to full-model finetuning, whereas ReFT methods (Wu et al., 2024a) operate on frozen model representations rather than weights, learning task-specific interventions that efficiently steer model behavior during inference.
@@
  • Reflection and refinement. Self-refine (Madaan et al., 2024) employs iterative generation, self-generated feedback, and refinement, enabling large language models to adjust based on feedback after each generative cycle. Reflexion (Shinn et al., 2023) expands on the ReAct framework by integrating an evaluator to assess action trajectories and utilizing an LLM to generate verbal self-reflections to provide feedback for future trials. CRITIC (Gou et al., 2023a) utilizes external tools, such as knowledge bases and search engines, to validate the actions produced by LLMs, then employs external knowledge for self-correction to minimize factual errors.
@@
  • Integrating language models, world models, and agent models. While language models fall short of consistent reasoning and planning in various scenarios, world and agent models can provide essential elements of human-like deliberative reasoning, including beliefs, goals, anticipation of consequences, and strategic planning. The LAW framework (Hu and Shu, 2023) suggests reasoning with world and agent models, with language models serving as the backend for implementing the system or its components. This framework combines three models in a cognitively grounded way, fostering more robust and versatile reasoning capabilities. Within this framework, Reasoning via Planning (RAP) (Hao et al., 2023) prompts an LLM to function as an agent model, guided by the same LLM acting as the world model, which predicts the next state of the reasoning after applying an action to the current state. BIP-ALM (Jin et al., 2024) and LIMP (Shi et al., 2024b) use language models as the planner in agent models, leading to an improved Theory of Mind capacity compared to using language models to infer other agents’ mental states directly. Recent studies have explored the potential for using language models to generate goals (Xie et al., 2023) or rewards (Yu et al., 2023; Kwon et al., 2023b; Ma et al., 2023a) in agent models to guide planning. Integrating these approaches, neural-symbolic methods can bridge the gap between the abstract reasoning facilitated by LLMs and the structured decision-making processes inherent in world and agent models. Logic-LM (Pan et al., 2023) apply symbolic execution on logical reasoning. Symbol-LLM (Xu et al., 2023b) unifies neural-symbolic applications under a Symbol+Delegation setting.
@@
  • Foundation models need to learn causation for better understanding and generalization. The foundation models rely heavily on patterns identified in their training data, which do not always capture the depth and breadth of human knowledge and experiences. Furthermore, these models often operate based on patterns extracted from data without truly comprehending the underlying causal relationships. Zečević et al. (2023) describe how LLMs might superficially replicate causal relationships but lack the underlying causal mechanisms, leading them to be more like “causal parrots” rather than genuinely causal models. Jin et al. (2023a) presents a challenging dataset for causal reasoning and suggests that LLMs may still be far from reasoning reliably about causality. Future advancements in AGI should focus on learning causation over correlation, thereby achieving better generalization and deeper understanding.
@@
  • AGI should get better at social reasoning to enhance interactions with humans and other agents. Current AI models lack a robust Theory of Mind, the ability to understand the mental states of others (Sap et al., 2022; Ullman, 2023; Jin et al., 2024; Shi et al., 2024b). Improving this capability is essential for AGI systems to safely and effectively interact with humans and other agents in an open-ended manner. Understanding social cues and norms is central to this development, as it allows AGIs to interpret and respond to implicit communication and behavioral expectations in varied contexts (Puig et al., 2020; Zhang et al., 2023b). Advancements in social reasoning in AGI systems could lead to more empathetic and context-aware technologies, ensuring that these systems engage harmoniously and meaningfully in human societies.
  • AGI should solve the challenge in explainability and transparency, thereby enhancing their reliability in decision-making. Currently, most AI systems lack these qualities, making it difficult to understand how they arrive at specific conclusions or answers (Chen et al., 2023k). Techniques that aim to elicit reasoning in natural language do not consistently align with the actual reasoning processes used by the models, and the explanations generated can be systematically misleading (Bowman, 2023). This limitation hinders their reasoning abilities and creates significant challenges when decision-making requires justification or auditing, particularly in fields like healthcare or law. Recent work on sparse autoencoders, specifically using them to address polysemanticity in neural networks, has shown promise in enhancing the explainability of AI models (Cunningham et al., 2023; Gao et al., 2024; Templeton, 2024). These autoencoders help in disentangling and isolating meaningful features within neural networks, leading to more interpretable and mono semantic features that are easier to understand. Incorporating similar techniques to interpret neural networks, and developing more techniques to explain different decision-making process of AGI systems can facilitate more trustworthy and accountable AI applications.
@@
  • Future AGI systems aim for dynamic reasoning across domains, ethical and efficient planning, and human-like intelligence at unparalleled scales and speeds. We are still far from achieving AGI-level capability that allows reasoning and planning across varied domains without retraining or human oversight (Saparov et al., 2024). The journey involves enhancing AI systems to transfer knowledge and skills across vastly different areas, enabling them to address unforeseen situations efficiently. A key development focus is creating algorithms that can plan at various levels of abstraction, from broad strategic goals to the specifics of detailed actions. Additionally, AI systems urgently need to become better at managing resources—such as time, energy, and costs—more efficiently during the planning phases. Equally critical is ensuring that these planning processes adhere to ethical standards and safety regulations, especially in sensitive sectors, to avoid misuse or unintended outcomes. Advancements in these areas will collectively move us closer to realizing AGI with robust, versatile planning capabilities.